The data below represent the number of days absent, x, and the final grade, y, for a sample of college students at a large university. Complete parts (a) through (e) below No. of absences, x 0 1 2 3 4 5 6 7 8 9 Final grade, y 88.9 86.0 82.9 80.3 77.4 72.9 63.4 67.7 64.7 61.7 (a) Find the least-squares regression line treating the number of absences, x, as the explanatory variable and the final grade, y, as the response variable.
Answer #1

(a)

Following table shows the calculations:

No. Of absences, X Final grade, Y X^2 Y^2 XY
0 88.9 0 7903.21 0
1 86 1 7396 86
2 82.9 4 6872.41 165.8
3 80.3 9 6448.09 240.9
4 77.4 16 5990.76 309.6
5 72.9 25 5314.41 364.5
6 63.4 36 4019.56 380.4
7 67.7 49 4583.29 473.9
8 64.7 64 4186.09 517.6
9 61.7 81 3806.89 555.3
Total 45 745.9 285 56520.71 3094

Sample size: n=10

Now,

(ソ)-= 884,029 S.

S_{xx}=\sum x^{2}-\frac{\left (\sum x \right )^{2}}{n}=82.5

S_{xy}=\sum xy-\frac{\left (\sum x \right )\left (\sum y \right )}{n}=-262.55

Slope of the regression equation is

b_{1}=\frac{S_{xy}}{S_{xx}}=-3.182424242\approx -3.18

and intercept of the equation will be

b_{0}=\frac{1}{n}(\sum y - b_{1} \sum x)=88.9

So the regression equation will be

y'=88.9-3.18x